Management accounting 1

Stilton Audio is a producer of speakers and amplifiers which are housed in plastic cabinets. Currently production of the plastic cabinets as well as final product assembly is completed in-house.

 

For accounting purposes, Stilton Audio uses the weighted average method, where plastic cabinets speaker and amplifier components are added at the beginning of the assembly process.

 

The following are cost and production data for the assembly process for April:

Production:

Units in process, April 1, 60 percent complete                60,000

Units completed and transferred out                                150,000

Units in process, April 30, 20 percent complete             30,000

Costs:

WIP, April 1:

Plastic cabinets                                                               $ 1,200,000

Speaker and amplifier components                              12,600,000

Conversion costs                                                               5,400,000

Costs added during April:

Plastic cabinets                                                                    2,400,000

Speaker and amplifier components                               25,200,000

Conversion costs                                                                 8,640,000

Required:

  1. Prepare a physical flow analysis for the assembly department for April.

  2. Calculate equivalent units of production for the assembly department for April.

  3. Calculate unit cost for the assembly department for April.

  4. Calculate the cost of units transferred out and the cost of EWIP inventory.

  5. Prepare a cost reconciliation for the assembly department for April.
    Answer #1

    1. Prepare a physical flow analysis for the assembly department for April.


    Physical units
    Physical units
    Beginning Units60000Units completed and transferred out150000
    Unit started120000Ending units30000
    Total units180000Total units180000

    2. Calculate equivalent units of production for the assembly department for April.


    PlastickSpeaker and amplifierConversion Cost
    Unit completed and transferred out150000150000150000
    Ending work in process30000300006000
    Equivalent unit180000180000156000








    3. Calculate unit cost for the assembly department for April.


    PlasticSpeaker and amplifier componentsConversion Cost
    Beginning work in process1200000126000005400000
    Cost added2400000252000008640000
    Total Cost36000003780000014040000
    Equivalent Unit180000180000156000
    Cost per equivalent unit2021090

    4. Calculate the cost of units transferred out and the cost of EWIP inventory.

    Cost of unit transferred out = (20+210+90)*150000 = $48000000

    Cost of ending work in process = (230*30000+90*6000) = 7440000

    5. Prepare a cost reconciliation for the assembly department for April.

    Cost to be account for
    Beginning work in process19200000
    Cost added36240000
    Total cost to be accounted for55440000
    Cost accounted as
    Cost of units transferred out48000000
    Cost of ending work in process7440000
    Total Cost accounted as55440000





    answered by: Tinnymove